How would you prepare a solution of 0.6mol L-1 HCl from 20.0mL of 4 mol L-1 stock HCl?
1
Expert's answer
2020-08-02T15:52:27-0400

Solution

Moles of HCl in first solution = moles of HCl after dilution

Moles = concentration * volume

Initial moles = 0.6* V

Finally moles = (20/1000*4)

0.6*V = 0.02*4

V = (.02*4)/0.6

= 8/60

V = 0.134litre

Or V= 134 ml. Approx.

So increased volume = 134 - 20

114 ml.

So we will have to add 114 ml. Of water for to prepare final solution.